Using a standard algorithm, calculate and show your working: 3- 1 2/3×5/6



1
Expert's answer
2021-05-23T19:15:02-0400

We calculate the given equation to get:

"3-\frac{12}{3}\:\times \frac{5}{6}=" "3-\frac{60}{18}"

"\frac{3}{1}-\frac{60}{18}=\frac{54-60}{18}" "=-\frac{6}{18}=-0.3333"